solve for c.
1945.4°=9/5c + 32

Answer by mananth(16946)

1945.4°=9/5c + 32
...
-32
1945.4-32=9/5 *c
1913.4=9/5*c
multiply by 5
5*1913.4=9c
/9
5*1913.4/9 =c
c=1063
...
m.ananth@hotmail.ca